Two spaceships are traveling, one in front of the other, in the same direction, at the same speed of 1.64 108 m/s relative to the earth. A laser pulse is emitted by one of the ships and is reflected back by a mirror on the other ship. The total round trip time for the laser pulse is 4.20 10-3 s. How long would it take for a small landing module to travel between the ships if its speed, relative to the ships, is 1.03 103 m/s?
Answer #1

As given the spaceships are travelling with same speed in same direction , So they stationary with respect to each other.

Hence The distance between then is given as : speed × time

d = s × t

Given, time to round trip = 4.20 × 10^-3 s

time taken to reach the mirror = 4.2/2 × 10^-3 = 2.10 × 10^-3 s

Speed of laser pulse = c = 3×10^8 m/s

Putting these value :

Distance (d) = 3×10^8 × 2.1×10^-3 = 6.3 × 10^5 m

Time taken by module to travel between the ships (T) = distance/speed

Given, Speed of module(s) = 1.03 × 10^3 m/s

Distance = 6.3×10^5 m

T = d/s = (6.3×10^5) / 1.03×10^3 = 6.1165 × 10^2 s

Time taken by landing module to travel between the ships is (T) = 611.65 second.

​​​​​